OpenStudy (anonymous):

If f(n) = - 5n - 2, then f(3) is what?

OpenStudy (anonymous):

plug 3 in for n

OpenStudy (angela210793):

f(3)=-5*3-2=-15-2=-17

OpenStudy (anonymous):

simplify (-5*3)-2. what do u get